A nurse is to give Fortaz to a child at an amount of 50.0 mg per kilogram of body weight. The child weighs 56.1 pounds. Fortaz is available in an oral suspension labeled 100 mg/mL. How many mL would the nurse give ? (1 kg = 2.2 pounds)
Answer #1

Ans:

Given:

Ordered recommended dose=50 mg /kg

Dose on hand=100 mg/ml

Weight of child=56.1 pounds

Solution:

Convert pounds to kg

Divide pounds  by 2.2

56.1÷2.2 = 25.5 kg

Therefore weight of child = 25.5 kg

Dose of child = weight × Recommended dose

= 25.5 × 50 = 1275 mg

If 1 ML contains 100 mg

Then x ML contains 1275 mg

Cross multiplication;

1 × 1275 = 100 × x

Therefore ;

X= 1275 ÷ 100 = 12.75 ml

Therefore nurse should give 12.75 ml Fortaz to the child.